OpenStudy (anonymous):

need help! If f(x) = x2 and g(x) = x + 6, find g(f(0)). A.-6 B.6 C.12 D.36

Nnesha (nnesha):

\[\large\rm g(\color{reD}{f(0)})\] first find f(0) substitute x for 0 in f(x) function then plug the result into tthe g(x) equation for x

Nnesha (nnesha):

make sense ?

OpenStudy (anonymous):

yes it do make sense thank you so much

Nnesha (nnesha):

alright let me know if you have a question :=))

OpenStudy (idealist10):

f(x)=x^2 g(x)=x+6 ----------- g(f(0))=g(0^2)=g(0)=0+6=6 Answer: 6

OpenStudy (anonymous):

thanks idealist10